For the metal salt indicated in the procedure, you
Answer #1

Molar mass is the combination atomic mass in gm present in 1 mole and unit is gm/mol

Ni(NO_3)_2.6H_2O

Mass of Ni = 58.69 gm

Mass N = 14.007 gm

Mass of O =15.999 gm

Mass of H = 1.007 gm

Now molar mass of Ni(NO_3)_2.6H_2O

1\times 58.69(Ni)+2\times 14.007(N)+12\times 15.999(O)+12\times 1.008(H)=290.786gm/mol

Here all atomic mass are given per mole and written in gm.